# Calibrating Bayesian Generative Machine Learning for Bayesiamplification

### v1

@@ Abstract
Recently, combinations of generative and Bayesian machine learning have been introduced in particle physics for both fast detector simulation and inference tasks. These neural networks aim to quantify the uncertainty on the generated distribution originating from limited training statistics. The interpretation of a distribution-wide uncertainty however remains ill-defined. We show a clear scheme for quantifying the calibration of Bayesian generative machine learning models. For a Continuous Normalizing Flow applied to a low-dimensional toy example, we evaluate the calibration of Bayesian uncertainties from either a mean-field Gaussian weight posterior, or Monte Carlo sampling network weights, to gauge their behaviour on unsteady distribution edges. Well calibrated uncertainties can then be used to roughly estimate the number of uncorrelated truth samples that are equivalent to the generated sample and clearly indicate data amplification for smooth features of the distribution.
@@ 3 Bayesian Continuous Normalizing Flows
Following [23], we first introduce the flow mapping $\phi_{t}:[0,1]\times\mathds{R}^{d}\rightarrow\mathds{R}^{d}$ parameterized by a time parameter $t\in[0,1]$. In analogy to the application of multiple blocks in a coupling-block flow [22], the change of the flow mapping between target and latent space is determined by an ordinary differential equation (ODE)
@@ 3.2 Markov Chain Monte Carlo
for a large range of noise parameter settings. Both the inverse temperature parameter $\lambda$ and the noise parameter $\sigma$ tune the predicted uncertainties. In theory small $\lambda$ and high $\sigma$ will result in high error estimates, albeit in practice the dependence on the inverse temperatures is very weak. We thus limit ourselves to adapting the noise parameter to align the generated uncertainties.
@@ 4.1 Gamma Function Ring
Similar to previous studies on data amplification [5], we employ the CNF on a low-dimensional ring distribution. Within this paper, we limit the study to two dimensions for illustrative purposes and to reduce computational costs. Nevertheless, the calibration can be executed analogously for higher dimensional distributions. We generate samples from a ring distribution with an unsteady edge at a radius of $r=4$, by sampling in spherical coordinates from
@@ 4.2 Hyperparameter Choices
Due to the low dimensionality of the toy example, we do not need to employ complicated architectures to obtain a good approximation of the vector-field $\tilde{v}_{t}(\cdot,\theta)$. Based on a small grid search, a Multi-Layer Perceptron with 3 layers of 32 nodes and ELU activation is deemed to be sufficient. Each of the 3 layers takes the time variable $t$ as an additional input. The neural network part of the CNF thus totals a mere $2498$ parameters.
When parameterizing the weight posterior approximation $\tilde{\pi}(\theta)$ as an uncorrelated Normal distribution, as is standard in VIB [17], the number of parameters consequently doubles. For VIB we train using the Adam optimizer [32] at a learning rate of $10^{-3}$ for up to $250$k epochs of $10$ batches of $1000$ datapoints each. To prevent overfitting, we evaluate the model at the earliest epoch after convergence of the KL-loss term. This point depends on the choice of $k$ and varies between $75$k for $k=50$ and $250$k for $k=1$. We do $5$ runs each for multiple values of $k\in[1,5,10,50]$ to regulate the uncertainty quantification.
For the AdamMCMC sampling, we start the chain from a pretrained model. The model is first optimized for $2500$ epochs (Adam with learning rate of $10^{-3}$) using only the CFM-loss (3.4). For the deterministic model, this is enough to converge. We then run the sampling at a the same learning rate as the optimization with $\sigma_{\Delta}\approx 50$ and $\lambda=1.0$. We add a sample to the collection at intervals of $100$ epochs, to ensure the independence of the sampled weights. To adjust the calibration, we scan the noise value at four points $\sigma\in[0.01,0.05,0.1,0.5]$. Once again, we calculate $5$ chains per noise parameter setting.
@@ 6 Bayesiamplification
We will now use the posterior mean and variance to construct an estimator $\hat{t}_{j}$ of the Poisson equivalent to the per-bin predictions. Using only the mean $\hat{t}_{j}\overset{!}{=}\bar{g}_{j}$, the equivalent will simply be the generated statistics. Thereby, we would disregard the correlations in the generated data through limited training data completely.
By instead equating the variance of the BNN to that of the equivalent uncorrelated set $\hat{t}_{j}~{}\overset{!}{=}~{}\sigma_{\bar{g}_{j}}$, we would introduce an unwanted dependence on the uncertainty prediction. Overestimated uncertainties would lead to an overestimation of the statistical power.
@@
$\frac{1}{\sqrt{\hat{t}_{j}}}\overset{!}{=}\frac{\sigma_{\bar{g}_{j}}}{\bar{g}_{j}}\quad\Longleftrightarrow\quad\hat{t}_{j}=\frac{\bar{g}_{j}^{2}}{\sigma_{\bar{g}_{j}}^{2}}.$ (6.1)
@@ Acknowledgments
Code
https://github.com/sbieringer/Bayesiamplify provides the code used in this analysis.

### v2

@@ Abstract
Recently, combinations of generative and Bayesian deep learning have been introduced in particle physics for both fast detector simulation and inference tasks. These neural networks aim to quantify the uncertainty on the generated distribution originating from limited training statistics. The interpretation of a distribution-wide uncertainty however remains ill-defined. We show a clear scheme for quantifying the calibration of Bayesian generative machine learning models. For a Continuous Normalizing Flow applied to a low-dimensional toy example, we evaluate the calibration of Bayesian uncertainties from either a mean-field Gaussian weight posterior, or Monte Carlo sampling network weights, to gauge their behaviour on unsteady distribution edges. Well calibrated uncertainties can then be used to roughly estimate the number of uncorrelated truth samples that are equivalent to the generated sample and clearly indicate data amplification for smooth features of the distribution.
@@ 3 Bayesian Continuous Normalizing Flows
Let $x\in\mathds{R}^{d}$ be a point in the data set $\mathcal{D}$. Following [30] , we first introduce the flow mapping $\phi_{t}:[0,1]\times\mathds{R}^{d}\rightarrow\mathds{R}^{d}$ parameterized by a time parameter $t\in[0,1]$. In analogy to the application of multiple blocks in a coupling-block flow [22], the change of the flow mapping between target and latent space is determined by an ordinary differential equation (ODE)
@@ 3.2 Markov Chain Monte Carlo
for a large range of noise parameter settings. If the added noise $\sigma$ is low, the results remain close to the stochastic optimization without error estimates close to zero, but if the noise levels are high, the random walk through parameter space dominates and the algorithm does not converge to a sensible parameter values. This behaviour is masked by diminishing acceptance probabilities for very low and very high $\sigma$ [21].
Both the inverse temperature parameter $\lambda$ and the noise parameter $\sigma$ tune the predicted uncertainties. In theory small $\lambda$ and high $\sigma$ will result in high error estimates, albeit in practice the dependence on the inverse temperatures is very weak. We thus limit ourselves to adapting the noise parameter to align the generated uncertainties.
@@ 4.1 Gamma Function Ring
Similar to previous studies on data amplification [5], we employ the CNF on a low-dimensional ring distribution. Generative Architectures often struggle with changes in the topology between latent space, typically Normal distributed, and data space [33]. The ring structure reflects this “topological worst case”. A generalization of the results from a similar, topologically complicated, but low-dimensional toy to high-dimensional, simulated, and topologically less problematic calorimeter images was performed in [6]. In this paper, we focus on the calibration of generative uncertainties and draw a connection to data amplification. We thus limit the study to two dimensions for illustrative purposes and to reduce computational costs. Nevertheless, the calibration can be executed analogously for higher dimensional distributions. We generate samples from a ring distribution with an unsteady edge at a radius of $r=4$, by sampling in spherical coordinates from
@@ 4.2 Hyperparameter Choices
Due to the low dimensionality of the toy example, we do not need to employ complicated architectures to obtain a good approximation of the vector-field $\tilde{v}_{t}(\cdot,\theta)$. Based on a small grid search, a Multi-Layer Perceptron with 3 layers of 32 nodes and ELU activation is sufficient to reproduce the training data well. Each of the 3 layers takes the time variable $t$ as an additional input. The neural network part of the CNF thus totals a mere $2498$ parameters.
When parameterizing the weight posterior approximation $\tilde{\pi}(\theta)$ as an uncorrelated Normal distribution, as is standard in VIB [17], the number of parameters consequently doubles. For VIB we train using the Adam optimizer [32] at a learning rate of $10^{-3}$ for up to $250$k epochs of $10$ batches of $1000$ datapoints each. To prevent overfitting, we evaluate the model at the earliest epoch after convergence of the KL-loss term. This point depends on the choice of $k$ and varies between $75$k for $k=50$ and $250$k for $k=1$. We do $5$ runs each for multiple values of $k\in[1,5,10,50]$ to regulate the uncertainty quantification. For this range of $k$ we have previously found sensible density estimation and optimization convergence trough performing a $\log$-linearly spaced scan in $k\in[10^{-4},10^{5}]$) with only one run per parameter choice.
For the AdamMCMC sampling, we start the chain from a pretrained model. The model is first optimized for $2500$ epochs (Adam with learning rate of $10^{-3}$) using only the CFM-loss (3.4). For the deterministic model, this is enough to converge. We then run the sampling at a the same learning rate as the optimization with $\sigma_{\Delta}\approx 50$ and $\lambda=1.0$. This choice of $\sigma_{\Delta}$ ensures high acceptance rates, while the choice of $\lambda$ reflects sampling from the untempered posterior distribution, as per (3.11). We add a sample to the collection at intervals of $100$ epochs, to ensure the independence of the sampled weights. To adjust the calibration, we scan the noise value at four points $\sigma\in[0.01,0.05,0.1,0.5]$. This parameter span is based on a $\log$-linearly spaced scan in $\sigma\in[10^{-4},10]$). Once again, we calculate $5$ chains per noise parameter setting.
@@ 6 Bayesiamplification
We will now use the posterior mean and variance to construct an estimator $\hat{t}_{j}$ of the Poisson equivalent to the per-bin predictions. Using only the mean $\hat{t}_{j}:=\bar{g}_{j}$, the equivalent will simply be the generated statistics. Thereby, we would disregard the correlations in the generated data through limited training data completely.
By instead equating the variance of the BNN to that of the equivalent uncorrelated set $\hat{t}_{j}~{}:=~{}\sigma_{\bar{g}_{j}}^{2}$, we would introduce an unwanted dependence on the uncertainty prediction. Overestimated uncertainties would lead to an overestimation of the statistical power.
@@
$\frac{1}{\sqrt{\hat{t}_{j}}}:=\frac{\sigma_{\bar{g}_{j}}}{\bar{g}_{j}}\quad\Longleftrightarrow\quad\hat{t}_{j}=\frac{\bar{g}_{j}^{2}}{\sigma_{\bar{g}_{j}}^{2}}.$ (6.1)
@@ Acknowledgments
Data and Code
https://github.com/sbieringer/Bayesiamplify provides the code for simulating the toy example and conducting this analysis.